40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 10 16:52:04 2020, gautam, Update, BHD, Workable setup prepared 
    Reply  Tue Aug 11 23:20:29 2020, gautam, Update, BHD, Some first tests with air BHD setup PRMI_RFlock.pdfPRMI_RFlock_fullscale.pdf
       Reply  Fri Aug 21 15:27:00 2020, gautam, Update, BHD, Better BHD mode-matching IMG_7548.JPGBHD_MM.pngPRMI_DCPDs.pdf
          Reply  Wed Aug 26 00:52:55 2020, gautam, Update, BHD, BHD activities 
Message ID: 15514     Entry time: Tue Aug 11 23:20:29 2020     In reply to: 15513     Reply to this: 15535
Author: gautam 
Type: Update 
Category: BHD 
Subject: Some first tests with air BHD setup 

Some tests done today:

All of these tests were done with the PRMI locked with carrier resonant in the recycling cavity (i.e. sidebands rejected to REFL port). I then actuated the BS length DOF with a sine wave at 311.1 Hz, 40 cts amplitude (corresponding to ~8 pm of peak-to-peak displacement).

  1. Attempt to balance the DCPDs
    • I tried to tune the digital gains of the two DCPDs so as to minimize the appearance of this line in the SUM channel
    • but no matter how I tuned the gains, I couldn't make the line in the SUM channel disappear entirely - in fact, the best I could do was to make the line height in SUM and NULL channels (yes I recognize the poor channel name choice, I'll change "NULL" to "DIFF" at the next model recompile) the same. See Attachment #1.
    • The lobes around the main peak are indicative of some scattering?
    • Attachment #2 shows a wider frequency range. The homodyne phase isn't controlled, so the "NULL" channel is not necessarily measuring the correct quadrature to be sensing MICH motion.
    • I think I can back out something about the contrast defect from this fact, but I need to go back to some modeling.
  2. A simple test of the homodyne phase actuator
    • I wanted to check that this PI S320 piezo actually allows me to actuate the optical path length of the local oscillator.
    • I'm using the OMC HV driver to drive said PZT - so there are two DAC channels available, one to dither the optic and one to apply a control signal. I think mainly this is to avoid using up DAC range for the dither signal, the overall dynamic range is still limited by the HV supply.
    • I can't find the maximum voltage that can be applied on the datasheet - so conservatively, I limited the HV output to saturate at 100 V DC, as this is the maximum for the S330 piezos used for green steering, for which there is a manual.
    • The S320 manual does say the full stroke of each PZT element is 10 um - so the actuation coefficient is ~100 nm/V. I then drove this actuator with a sine wave of 500 cts amplitude, at 314.1 Hz (corresponding to 15 nm of motion). With only the LO beam incident on the PDs, I saw no signal in either DCPD - as expected, so this was good.
    • Then, with the PRMI locked, I repeated the test. If there is no DC light field (as expected for the PRMI in this configuration), I wouldn't expect this drive signal to show up in the DCPDs. But in fact, I do. Again, this supports the presence of some (for now unquantified) contrast defect.

While it would seem from these graphs that the RIN of the LO beam at these frequencies is rather high, it is because of the ADC noise. More whitening (to be installed in the coming days) will allow us to get a better estimate, should be ~1e-6 I think.

I was just playing today, still need to setup some more screens, DTT templates etc to do more tests in a convenient way.

Now, I can think about how to commission this setup interferometrically.

Attachment 1: PRMI_RFlock.pdf  15 kB  Uploaded Wed Aug 12 00:53:20 2020  | Hide | Hide all
PRMI_RFlock.pdf
Attachment 2: PRMI_RFlock_fullscale.pdf  104 kB  Uploaded Wed Aug 12 00:59:29 2020  | Hide | Hide all
PRMI_RFlock_fullscale.pdf
ELOG V3.1.3-